# LATex: Leveraging Attribute-based Text Knowledge for Aerial-Ground Person Re-Identification

### v1

@@ 4.2 Prompted Attribute Classifier Group
To enhance attribute perception, we propose the Prompted Attribute Classifier Group (PACG), which integrates attribute label information and exploits interdependencies among attributes. More specifically, we define PACt as the classifier for the $t$-th attribute. Then, the attribute features $F_{a}^{t}$ is defined as the concatenation of the global feature $F_{v}^{L}$ and corresponding attribute prompt $\tilde{P}^{L,t}_{v}$, where $F_{v}^{L,t}\in\mathbb{R}^{C}$ is the $t$-th attribute prompts of $F_{v}^{L}$:
$F_{a}^{t}=[F_{v}^{L},\tilde{P}^{L,t}_{v}].$ (5)
@@
To learn more discriminative attribute-aware features, we utilize the interdependencies of different attributes. Formally, we denote other attribute prompts as $\hat{P}^{L}_{v}$, i.e., $\hat{P}^{L}_{v}=\{P_{v}^{L,1},\cdots,P_{v}^{L,t-1},P_{v}^{L,t+1},\cdots,P_{v}^{L,T}\}$. Then, we can obtain interacted features as follows:
@@
where $\Theta(\cdot)$ is the multi-head cross attention (Vaswani et al., 2017). $Q\in\mathbb{R}^{C}$, $K\in\mathbb{R}^{C}$ and $V\in\mathbb{R}^{C}$ are generated by the corresponding projection matrix $W_{q}$, $W_{k}$ and $W_{v}$, respectively. $\delta$ is the Softmax function. The residual structure enables the model to process input features with varying emphases more flexibly  (He et al., 2016). As observed in daily life, certain person attributes exhibit strong correlations (e.g., gender and clothing), while others show weaker correlations (e.g., height and weight) or are nearly independent (e.g., ethnicity and gender). Thus, we employ a residual-based Feed-Forward Network (FFN)  (Vaswani et al., 2017) to handle features obtained from two perspectives: direct prediction and attribute dependency-based prediction. This design allows the FFN to adapt to diverse scenarios by effectively capturing attribute correlations, if they exist, while avoiding excessive noise and additional computational overhead caused by irrelevant attribute pairs:
@@
where $\Phi(\cdot)$ represents the FFN and $\mathcal{O}_{t}$ is the final features used to predict attribute confidences. In addition, to align the visual and textual representations in different feature spaces, we transform the visual embedding $F_{a}^{t}$ and obtain attribute-based text tokens as follows:
@@
where $\Psi$ is a Multi-Layer Perceptron (MLP). The above attribute-based text tokens can be taken as the input of CLIP’s text encoder to improve the feature discrimination.
4.3 Coupled Prompt Template
Recently, visual-language models, such as CLIP have delivered outstanding performances in many computer vision and natural language processing tasks. As a fundamental component, text templates play an important role in text-missing tasks, such as person ReID. For example, Li et al. (Li et al., 2023) utilize identity-specific tokens to form a text template, i.e., “A photo of a [learnable tokens] person.” However, this kind of templates only focus on person appearances, ignoring the helpful attribute information.
To address the above issue, we propose a Coupled Prompt Template (CPT), which is presented as “A [view token] view photo of a [share tokens] [attribute tokens] person.” This template not only couples identity-independent and attribute-aware information, but also leverages comprehensive knowledge of visual-language models. Specially, the view token $V$ is an instance-level text token, which depends on the view of images. When the image is captured by a ground camera, the view token is “CCTV”. When the image is from an aerial camera, it is assigned to “UAV”. In addition, we formulate share tokens as “$[M_{1},M_{2},\cdots,M_{K}]$”, where $K$ is the total number of tokens. These instance-shared tokens serve as register tokens (Darcet et al., 2023) to enhance semantic feature representations. As for attribute tokens, we denote them as “$[A_{1},A_{2},\cdots,A_{T}]$”, which are obtained by PACG and have rich attribute information. By effectively utilizing these diverse tokens to form the structured sentence $\mathcal{S}$, our framework can fully leverage useful information from the attributes. Finally, we feed $\mathcal{S}$ to the text encoder $\mathcal{T}(\cdot)$ of CLIP to obtain the text feature $F_{d}$:
@@ 4.3 Coupled Prompt Template
To improve the feature representation ability, we concatenate $F_{d}$ and $F_{v}^{L}$ for person retrieval. With the CPT, our LATex can fully leverage the viewpoint invariance of person attributes to enhance the identify-related features.
4.4 Loss Functions
As illustrated in Fig. 3, we employ multiple loss functions to optimize our framework. For features obtained by AIE and CPT, we supervise them by the label smoothing cross-entropy loss (Szegedy et al., 2016) and triplet loss (Hermans et al., 2017):
@@ 4.4 Loss Functions
5 Experiment
5.1 Datasets and Evaluation Protocols
Datasets. We evaluate our methods on three AG-ReID benchmarks. AG-ReID.v1 (Nguyen et al., 2023a) is a challenging dataset, consisting of 21,983 images captured by ground and aerial cameras from 388 identities with 15 attributes. CARGO (Zhang et al., 2024b) is a large-scale synthetic dataset, consisting of 108,563 images captured by five aerial cameras and eight ground cameras from 5,000 identities. Meanwhile, AG-ReID.v1 contains two protocols for evaluation, i.e., A→G and G→A. CARGO contains four evaluation protocols, two of which are view-heterogeneous, while the other two are view-homogeneous. Finally, AG-ReID.v2 (Nguyen et al., 2024) is the extended version of AG-ReID.v1, incorporating three views: aerial (A), ground (G), and wearable device (W). Accordingly, the evaluations are expanded to include cross-view retrieval settings between A and W. Specifically, AG-ReID.v2 consists of 100,502 images from 1,615 unique identities.
Metrics. Following previous works, we employ the mean Average Precision (mAP) (Zheng et al., 2015) and Cumulative Matching Characteristic (CMC) (Moon & Phillips, 2001) at Rank-1 as evaluation metrics.
5.2 Implementation Details
Our proposed framework is implemented with PyTorch on one NVIDIA A100 GPU. We use the pre-trained CLIP, specially CLIP-Base-16, as our backbone. All input images are resized to $256\times 128$. To enhance the generalization ability, we utilize multiple augmentation techniques such as the random horizontal flipping, padding and random erasing (Zhong et al., 2020) for all inputs. While the model is training, the size of each mini-batch is 128 consisted of 16 identities and 8 instances of each identity. We fine-tune the model by the Adam optimizer with a base learning rate of 0.00035. The total epoch is 120. For the hyper-parameters, we set $\lambda_{1}$ and $\lambda_{2}$ to 0.25, and 1.0, respectively.
Table 1: Performance comparison with existing methods on AG-ReID.v1. The superscript symbol † indicates results based on full fine-tuning strategies. The best and second-best results are highlighted in bold and underline.
@@ 5.2 Implementation Details
OSNet(Zhou et al., 2021) 72.59 58.32 74.22 60.99
BoT(Luo et al., 2019) 70.01 55.47 71.20 58.83
SBS(He et al., 2023) 73.54 59.77 73.70 62.27
VV(Kumar et al., 2019, 2020) 77.22 67.23 79.73 69.83
ViT(Dosovitskiy et al., 2020) 81.28 72.38 82.64 73.35
Explain(Nguyen et al., 2023a) 81.47 72.61 82.85 73.39
VDT(Zhang et al., 2024b) 82.91 74.44 86.59 78.57
@@
Table 2: Performance comparison on CARGO under view-heterogeneous and view-homogeneous protocols. The superscript symbol † indicates results based on full fine-tuning strategies. The best and second-best results are highlighted in bold and underline. Performance is shown for view-heterogeneous protocols in red, and for view-homogeneous in blue.
@@
SBS(He et al., 2023) 50.32 43.09 31.25 29.00 72.31 62.99 67.50 49.73
PCB(Sun et al., 2021) 51.00 44.50 34.40 30.40 74.10 67.60 55.00 44.60
BoT(Luo et al., 2019) 54.81 46.49 36.25 32.56 77.68 66.47 65.00 49.79
MGN(Wang et al., 2018) 54.81 49.08 31.87 33.47 83.93 71.05 65.00 52.96
VV(Kumar et al., 2019, 2020) 45.83 38.84 31.25 29.00 72.31 62.99 67.50 49.73
AGW(Ye et al., 2021) 60.26 53.44 43.57 40.90 81.25 71.66 67.50 56.48
ViT(Dosovitskiy et al., 2020) 61.54 53.54 43.13 40.11 82.14 71.34 80.00 64.47
VDT(Zhang et al., 2024b) 64.10 55.20 48.12 42.76 82.14 71.59 82.50 66.83
@@
Table 3: Performance comparison on AG-ReID.v2. The superscript symbol † indicates results based on full fine-tuning strategies. The best and second-best results are highlighted in bold and underline.
Method A→C A→W C→A W→A
Rank-1 mAP Rank-1 mAP Rank-1 mAP Rank-1 mAP
Swin(Liu et al., 2021b) 68.76 57.66 68.49 56.15 68.80 57.70 64.40 53.90
HRNet-18(Wang et al., 2020) 75.21 65.07 76.26 66.17 76.25 66.16 76.25 66.17
SwinV2(Liu et al., 2022) 76.44 66.09 80.08 69.09 77.11 62.14 74.53 65.61
MGN(R50)(Wang et al., 2018) 82.09 70.17 88.14 78.66 84.21 72.41 84.06 73.73
BoT(R50)(Luo et al., 2019) 80.73 71.49 86.06 75.98 79.46 69.67 82.69 72.41
BoT(R50)+Attributes 81.43 72.19 86.66 76.68 80.15 70.37 83.29 73.11
SBS(R50)(He et al., 2023) 81.96 72.04 88.14 78.94 84.10 73.89 84.66 75.01
SBS(R50)+Attributes 82.56 72.74 88.74 79.64 84.80 74.59 85.26 75.71
ViT(Dosovitskiy et al., 2020) 85.40 77.03 89.77 80.48 84.65 75.90 84.27 76.59
V2E(ViT)(Nguyen et al., 2024) 88.77 80.72 93.62 84.85 87.86 78.51 88.61 80.11
LATex 87.18 79.92 90.09 83.50 85.86 79.07 87.52 80.93
LATex† 89.13 83.50 91.35 86.35 89.01 82.85 89.32 83.30
Table 4: Training cost comparison of different methods.
@@
GPU Memory(G) 0.075 0.078 0.079
Table 5: Ablation results of key modules.
@@
Table 6: Performance comparison with different features.
@@
5.3 Performance Comparison
We compare our proposed method with existing methods on three AG-ReID benchmarks in Tab. 1, Tab. 2 and Tab. 3.
LATex. On AGReID.v1, our LATex achieves a Rank-1 accuracy of 84.41% and an mAP of 75.85% under the evaluation protocol A→G. Compared with VDT, it delivers 1.5% and 1.41% improvements in Rank-1 and mAP, respectively. As for the evaluation protocol G→A, the performance of our LATex is 88.88% Rank-1 and 79.19% mAP, showing a similar trade in evaluation protocol A→G. The stable improvements on two evaluation protocols clearly demonstrate the importance of further exploring attribute information in AG-ReID. On two large-scale datasets, CARGO and AG-ReID.v2, our LATex achieves highly competitive performance. It is worth noting that, since there are no attribute annotations on CARGO, we remove PACG. Our model achieves superior performance with significantly fewer trainable parameters, fully validating its effectiveness.
LATex†. We believe the limited trainable parameters constrain the performance of LATex. To enable a fairer comparison with previous methods based on full fine-tuning backbones, we introduce a LATex variant, LATex†, which adopts the same strategy. Compared to LATex, LATex† achieves significant performance improvements across all benchmarks. For example, on the CARGO dataset under the ALL protocol, LATex surpasses VDT with a Rank-1 accuracy of 66.99%, while LATex† further boosts this metric to 76.96%, achieving an almost 10% increase. This fully demonstrates the scalability of our methods.
[…]
5.4 Training Cost Comparison
In Tab. 4, we present a comprehensive comparison of costs on AG-ReID.v1. Compared with ViT and VDT, our LATex only utilizes about 40% of the trainable parameters while introducing minimal overhead.
5.5 Ablation Studies
To demonstrate the effectiveness of our proposed modules, we evaluate them on the AG-ReID.v1 dataset. The result is shown in Tab. 5. Furthermore, we conduct comprehensive and systematic evaluations to demonstrate the effectiveness of the proposed LATex framework.
Effect of Different Modules. In Tab. 5, Model A is our baseline. It achieves a Rank-1 of 81.69% and an mAP of 72.36% under the protocol A→G, while attaining a Rank-1 of 83.89% and an mAP of 74.78% under the protocol G→A. With PACG, Model B increases the performance by 3.53% Rank-1 and 3.84% mAP scores under the protocol G→A. Obviously, its performance is already better than previous methods, such as VDT. With CPT, Model C achieves the best result across all evaluation metrics. These results clearly demonstrate the effectiveness of our key modules.
Effect of Leveraging Person Attributes. We further validate the effect of leveraging person attributes. As shown in Tab. 6, Model A and Model B are models without CPT. Model C and Model D are completed models. The key differences are the features used for retrieval. Here, $F_{a}$ is the concatenation of all $F_{a}^{t}$. As can be observed, the performance of Model B is superior to that of Model A. The performances of Model B and Model D are very comparable. The performance of Model C is the best. These results clearly highlight the effectiveness of leveraging attribute-based text knowledge for AG-ReID.
Role of Backbones. We emphasize that the core idea of our LATex is to extract person attributes, embed them as pseudo-texts, and feed them into a text encoder to achieve robust person ReID. With the strong vision-language alignment, we adopt CLIP as the backbone. Considering that our approach uses a different backbone from prior methods, we evaluate the effect of various backbones and report the performance on AG-ReID.v1 in Tab. 7. As can be seen, CLIP-based VDT and ViT-based LATex show a significant performance drop. The main reason is that they cannot fully leverage CLIP’s vision-language knowledge. These results clearly demonstrate that our LATex’s performance stems from effectively utilizing CLIP’s vision-language knowledge, rather than its inherent encoding capabilities.
Table 7: Performance comparison with different backbones.
Methods A→G G→A
@@ 5.5 Ablation Studies
Impact of Trainable Prompts. Fig. 4 shows the effect of the number of learnable prompts $P^{i}_{v}$. With the increase of trainable parameters, the performance of LATex remains stable, demonstrating its robustness to this factor.

### v2

@@ III-C Prompted Attribute Classifier Group
To explicitly predict person attributes using image information, we propose the Prompted Attribute Classifier Group (PACG), which integrates attribute information and exploits interdependencies among attributes. More specifically, we define PACt as the classifier for the $t$-th attribute. Then, the attribute feature $F_{a}^{t}$ is defined as the concatenation of the global feature $F_{v}^{L}$ and the corresponding attribute prompt $\tilde{P}^{L,t}_{v}$:
$F_{a}^{t}=[F_{v}^{L},\tilde{P}^{L,t}_{v}],$ (5)
@@
where $\tilde{P}^{L,t}\in\mathbb{R}^{C}$ is the $t$-th attribute prompt of $\tilde{P}_{v}^{L}$. To further enhance the attribute-aware features, we utilize the interdependencies of different person attributes. Formally, we denote other attribute prompts as $\hat{P}^{L}_{v}=\{P_{v}^{L,1},\cdots,P_{v}^{L,t-1},P_{v}^{L,t+1},\cdots,P_{v}^{L,T}\}$. Then, we can obtain interacted feature as follows:
@@
where $\Theta(\cdot)$ is the multi-head cross attention [59]. $Q\in\mathbb{R}^{C}$, $K\in\mathbb{R}^{C}$ and $V\in\mathbb{R}^{C}$ are generated by the corresponding projection matrix $W_{q}$, $W_{k}$ and $W_{v}$, respectively. $\delta$ is the Softmax function. The residual structure enables the model to process input features more flexibly [60]. As observed in daily life, certain person attributes exhibit strong correlations (e.g., gender and clothing), while others show weaker correlations (e.g., height and weight) or are nearly independent (e.g., ethnicity and gender). Thus, we employ a residual-based Feed-Forward Network (FFN) [59] to handle features obtained from two perspectives: direct prediction and attribute dependency-based prediction. This design allows the FFN to adapt to diverse scenarios by effectively capturing attribute correlations, if they exist, while avoiding excessive noises and additional computational overhead caused by irrelevant attribute pairs:
@@
where $\Phi(\cdot)$ represents the FFN and $\mathcal{O}_{t}$ is the final features used to predict attribute confidences. In addition, to align the visual and textual representations in different feature spaces, we transform the visual embedding $F_{a}^{t}$ and obtain attribute-based text representations as follows:
@@
where $\Psi$ is a Multi-Layer Perceptron (MLP). The above attribute-based text representations are then input into CLIP’s text encoder to enhance feature discrimination.
III-D Coupled Prompt Template
Recently, large-scale vision-language models [61], such as CLIP have delivered outstanding performance in many computer vision and natural language processing tasks. As a fundamental component, text templates play an important role in text-based person ReID. For example, Li et al. [20] utilize identity-specific prompt tokens to form a text template, i.e., “A photo of a [learnable tokens] person.” However, this kind of templates learn person attributes implicitly, lacking the explicit supervision and ignoring helpful view information.
To address the above issues, we propose a Coupled Prompt Template (CPT), which is presented as “A [view token] view photo of a [shared tokens] [attribute tokens] person.” This template not only couples identity-independent and attribute-aware information, but also leverages comprehensive knowledge of visual-language models. More specifically, the view token $V$ is an instance-level text token, which depends on the view of person images. When the image is captured by a ground camera, the view token is “CCTV”. When the image is from an aerial camera, it is assigned to “UAV”. In addition, we formulate shared tokens as $M_{1},M_{2},\cdots,M_{K}]$”, where $K$ is the total number of tokens. These instance-shared tokens serve as register tokens [62] to enhance semantic feature representations. As for attribute tokens, we denote them as $A_{1},A_{2},\cdots,A_{T}]$”, which are obtained by PACG and have rich attribute information. By effectively utilizing these diverse tokens to form the structured sentence $\mathcal{S}$, our framework can fully leverage the useful information from person attributes. Finally, we feed the sentence $\mathcal{S}$ to the text encoder $\mathcal{T}(\cdot)$ of CLIP to obtain the text feature $F_{d}\in\mathbb{R}^{C}$:
@@ III-D Coupled Prompt Template
To improve the feature discrimination, we concatenate $F_{d}$ and $F_{v}^{L}$ for person retrieval. With the CPT, our LATex can fully leverage the viewpoint invariance of person attributes to enhance the identity-related features.
III-E Loss Functions
As illustrated in Fig. 2, we employ multiple loss functions to optimize our framework. For features obtained by AIE and CPT, we supervise them by the label smoothing cross-entropy loss [63] and triplet loss [64]:
@@ III-E Loss Functions
IV Experiment
IV-A Datasets and Evaluation Metrics
Datasets. We evaluate our methods on three AG-ReID benchmarks. AG-ReID.v1 [9] is a challenging dataset, consisting of 21,983 images captured by ground and aerial cameras of 388 identities, each annotated with 15 attributes. Meanwhile, AG-ReID.v1 contains two protocols for evaluation, i.e., A→G and G→A. AG-ReID.v2 [49] is an extended version of AG-ReID.v1, incorporating three views: aerial (A), ground (G), and wearable device (W). Accordingly, the evaluations are expanded to include cross-view settings between A and W. Specifically, AG-ReID.v2 consists of 100,502 images from 1,615 unique identities. Finally, CARGO [11] is a large-scale synthetic dataset, consisting of 108,563 images captured by five aerial cameras and eight ground cameras from 5,000 identities. It is worth noting that there are no attribute annotations in this dataset. For its protocols, CARGO contains four evaluation protocols, two of which are view-heterogeneous, while the other two are view-homogeneous.
Metrics. Following previous works, we employ the mean Average Precision (mAP) [28] and Cumulative Matching Characteristic (CMC) [65] at Rank-1 as evaluation metrics.
IV-B Implementation Details
Our proposed framework is implemented with PyTorch on one NVIDIA A100 GPU. We use the pre-trained CLIP-Base-16 as our backbone. All input images are resized to $256\times 128$. To enhance the generalization ability, we utilize multiple augmentation techniques such as random horizontal flipping, padding and random erasing [66] for all inputs. While the model is training, the mini-batch size is 128 consisted of 16 identities and 8 instances of each identity. We fine-tune the model with Adam optimizer [67] with a base learning rate of 3.5e-4. A learning rate scheduling strategy is employed, combining a warm-up phase with cosine decay and a scaling factor of 0.01. The total epoch is 120. For the hyper-parameters, we set $\lambda_{1}$ and $\lambda_{2}$ in Eq. 11 to 0.25 and 1.0, respectively.
LATex†. LATex† is a variant of LATex that adopts the full fine-tuning strategy to ensure a fair comparison with other full fine-tuning AG-ReID methods. More specifically, we unfreeze all trainable parameters of the pre-trained CLIP vision and text encoders and update them with a learning rate of 5e-6. Other experimental settings, including the learning rate decay strategy, optimizer, and the number of training epochs, are kept consistent with LATex.
TABLE I: Performance comparison with different methods on AG-ReID.v1. The symbol † indicates results based on full fine-tuning strategies. The best and second-best results are highlighted in bold and underline, respectively.
@@ IV-B Implementation Details
OSNet[68] 72.59 58.32 74.22 60.99
BoT[39] 70.01 55.47 71.20 58.83
SBS[69] 73.54 59.77 73.70 62.27
VV[70] 77.22 67.23 79.73 69.83
ViT[71] 81.28 72.38 82.64 73.35
TransReID[41] 81.80 73.10 83.40 74.60
PFD[72] 82.30 73.60 82.50 73.90
PHA[73] 79.30 71.30 81.10 72.10
FusionReID[43] 80.40 71.40 82.40 74.20
CLIP-ReID[20] 79.44 70.55 84.20 73.05
PCL-CLIP[55] 82.16 73.11 86.90 76.28
Explain[9] 81.47 72.61 82.85 73.39
VDT[11] 82.91 74.44 86.59 78.57
DTST[53] 83.48 74.51 84.72 76.05
SeCap[52] 84.03 76.16 87.01 78.34
@@
TABLE II: Performance comparison on AG-ReID.v2. The superscript symbol † indicates results based on full fine-tuning strategies. The best and second-best results are highlighted in bold and underline, respectively.
Method A→C A→W C→A W→A
Rank-1 mAP Rank-1 mAP Rank-1 mAP Rank-1 mAP
Swin[74] 68.76 57.66 68.49 56.15 68.80 57.70 64.40 53.90
HRNet-18[75] 75.21 65.07 76.26 66.17 76.25 66.16 76.25 66.17
SwinV2[76] 76.44 66.09 80.08 69.09 77.11 62.14 74.53 65.61
MGN(R50)[77] 82.09 70.17 88.14 78.66 84.21 72.41 84.06 73.73
BoT(R50)[39] 80.73 71.49 86.06 75.98 79.46 69.67 82.69 72.41
BoT(R50)+Attributes 81.43 72.19 86.66 76.68 80.15 70.37 83.29 73.11
SBS(R50)[69] 81.96 72.04 88.14 78.94 84.10 73.89 84.66 75.01
SBS(R50)+Attributes 82.56 72.74 88.74 79.64 84.80 74.59 85.26 75.71
BoT(ViT)[39] 85.40 77.03 89.77 80.48 84.65 75.90 84.65 75.90
ViT[71] 85.40 77.03 89.77 80.48 84.65 75.90 84.27 76.59
TransReID[41] 88.00 81.40 90.40 84.50 87.60 80.10 87.70 81.10
FusionReID[43] 86.70 80.70 89.70 84.20 87.90 80.00 86.50 80.90
CLIP-ReID[20] 85.36 79.79 89.14 84.23 85.64 79.08 86.50 79.55
PCL-CLIP[55] 79.80 72.20 87.14 77.70 81.12 72.40 84.19 73.89
Explain[10] 87.70 79.00 93.67 83.14 87.35 78.24 87.73 79.08
VDT[11] 86.46 79.13 90.00 82.21 86.14 78.12 85.26 78.52
V2E(ViT)[49] 88.77 80.72 93.62 84.85 87.86 78.51 88.61 80.11
SeCap[52] 88.12 80.84 91.44 84.01 88.24 79.99 87.56 80.15
LATex 87.18 79.92 90.09 83.50 85.86 79.07 87.52 80.93
LATex† 89.13 83.50 91.35 86.35 89.01 82.85 89.32 83.30
TABLE III: Performance comparison on CARGO. The superscript symbol † indicates results based on full fine-tuning strategies. The best and second-best results are highlighted in bold and underline, respectively. Performance is shown for view-heterogeneous protocols in red, and for view-homogeneous in blue.
@@
SBS[69] 50.32 43.09 31.25 29.00 72.31 62.99 67.50 49.73
PCB[78] 51.00 44.50 34.40 30.40 74.10 67.60 55.00 44.60
BoT[39] 54.81 46.49 36.25 32.56 77.68 66.47 65.00 49.79
MGN[77] 54.81 49.08 31.87 33.47 83.93 71.05 65.00 52.96
VV[70] 45.83 38.84 31.25 29.00 72.31 62.99 67.50 49.73
AGW[79] 60.26 53.44 43.57 40.90 81.25 71.66 67.50 56.48
ViT[71] 61.54 53.54 43.13 40.11 82.14 71.34 80.00 64.47
TransReID[41] 73.70 64.70 64.40 55.90 85.70 77.90 85.00 71.80
FusionReID[43] 67.90 61.50 48.30 53.10 85.70 79.40 80.00 69.30
CLIP-ReID[20] 68.27 64.25 55.62 53.83 84.82 80.80 75.00 65.42
PCL-CLIP[55] 67.31 60.93 54.37 51.43 84.82 76.00 70.00 60.75
VDT[11] 64.10 55.20 48.12 42.76 82.14 71.59 82.50 66.83
DTST[53] 64.42 55.73 50.63 43.39 78.57 72.40 80.00 63.31
SeCap[52] 68.59 60.19 69.43 58.94 86.61 75.42 80.00 68.08
@@
TABLE IV: Training and inference cost comparison of different methods.
@@
TABLE V: Ablation results of key modules.
@@
TABLE VI: Performance comparison with different features.
@@
TABLE VII: Performance comparison with different backbones.
Method A→G G→A
@@
IV-C Performance Comparison
We compare our proposed method with other person ReID methods on three AG-ReID benchmarks in Tab. I, Tab. II and Tab. III. Experiments on these AG-ReID benchmarks clearly show impressive performance of our proposed method.
On the AG-ReID.v1, our LATex achieves a Rank-1 accuracy of 88.88% and an mAP of 79.19% under the evaluation protocol G→A. Compared with SeCap, our LATex represents improvements of 1.87% in Rank-1 and and 0.85% in mAP, respectively. As for the evaluation protocol A→G, the performance of our LATex is 84.41% Rank-1 and 75.85% mAP, showing very competitive results. The consistent improvements on two evaluation protocols clearly demonstrate the importance of leveraging attribute information in AG-ReID. On two large-scale datasets, CARGO and AG-ReID.v2, our LATex achieves highly competitive performance. It is worth noting that, since there are no attribute annotations on CARGO, we remove PACG. As a result, the final outputs of AIE, i.e., $\tilde{P}^{L}$, are directly passed to CPT as attribute representations. Our LATex achieves superior performance with fewer trainable parameters, fully validating its effectiveness.
To enable a fairer comparison with previous methods based on full fine-tuning, we introduce a LATex variant, namely LATex†. Compared with LATex, LATex† achieves significant performance improvements across all benchmarks. For example, on the CARGO dataset under all protocols, LATex surpasses DTST with a Rank-1 accuracy of 66.99%, while LATex† further boosts this evaluation metric to 76.96%, achieving an almost 10% increase. This fully demonstrates the scalability of our methods on powerful backbones.
[…]
CLIP-based ReID. Recently, CLIP has been used as a visual backbone for AG-ReID tasks. To keep the advance, we compare our proposed method with some typical CLIP-based ReID methods in Tab. I, Tab. II and Tab. III. All models are trained with the full fine-tuning strategy. We focus on this kind of comparisons on the CARGO benchmark, since CARGO does not comprise attribute annotations so that we can exclude the impact of additional information. As shown in Tab. III, though these methods show impressive performance in view-homogeneous scenarios, they fail to handle the drastic viewpoint changes. Consequently, they degrade significantly on CARGO. In contrast, our LATex is able to address this issue, thus performs well in both view-homogeneous and view-heterogeneous settings.
IV-D Training and Inference Cost Comparison
In Tab. IV, we provide a comprehensive comparison of the training and inference costs on AG-ReID.v1. During training, our model significantly reduces the trainable parameters to about 40% compared with VDT. As a result, when all other conditions remain the same, the resource requirements for training the model (e.g., the computational overhead of parameter update) can be reduced. As for the inference cost, LATex obtains notable performance gains in the cost of introducing minimal computational overhead compared with other AG-ReID methods. The inference speed is also very comparable.
(a) A→G protocol
(b) A→G protocol